Assume that the paired data came from a population that is normally distributed. Using a

0.05

significance level and

d equals=x minus??y,

find

d overbard?,

s Subscript dsd?,

the t test? statistic, and the critical values to test the claim that

mu Subscript d?dequals=0.

x

7

1

6

9

6

10

6

11

y

5

6

8

14

8

12

8

12

d overbardequals=

?(Round to three decimal places as? needed.)

s Subscript dsdequals=

?(Round to three decimal places as? needed.)

t equals=

?(Round to three decimal places as? needed.)

t Subscript alpha divided by 2t?/2equals=plus or minus±nothing

?(Round to three decimal places as? needed.)
Answer #1

Ans:

x 7 1 6 9 6 10 6 11
y 5 6 8 14 8 12 8 12
difference(d) 2 -5 -2 -5 -2 -2 -2 -1
mean= -2.125
std. dev= 2.232

From above data:

d--2.125

sd 2.232

Test statistic:

t=(-2.125-0)/(2.232/sqrt(8))

t=-2.693

df=8-1=7

critical t value(two tailed)=tinv(0.05,7)=+/-2.365